# ReferencesSuppReferences

### v3

@@ 4.1 Response of Interest Rates to Monetary Policy News
For a larger value of $\rho$ it takes larger values of $\delta$ to decrease the MAE and MSE of $\widehat{\beta}_{\mathrm{ES}}$ to those of $\widehat{\beta}_{\mathrm{oracle}}$. For example, for $\rho=0.5$ increasing the variance of the policy variable in the treatment sample nine-fold is not sufficient, while increasing it roughly thirty three times is. Similar features are found when the dependent variable is calibrated to the 5-Year real yield.
The bias of $\widehat{\beta}_{\mathrm{oracle}}$ is always smaller than that of $\widehat{\beta}_{\mathrm{ES}}$. This follows because for $\widehat{\beta}_{\mathrm{oracle}}$ positive and negative deviations from the true value $\beta$ cancel each other across simulation replications, while for $\widehat{\beta}_{\mathrm{ES}}$ these deviations tend to be positive since $\rho>0$. On the other hand, these deviations for $\widehat{\beta}_{\mathrm{ES}}$ tend to be smaller in magnitude, making its variance smaller than that of $\widehat{\beta}_{\mathrm{oracle}}$ for large values of delta. Focusing only on bias does not account for the higher precision of $\widehat{\beta}_{\mathrm{ES}}$. A user concerned about both accuracy and precision should rather focus on performance measures such as MSE or MAE, which incorporate both bias and variance.
@@
Values of $\rho=0.50$ and 0.40 represent quite strong empirical contemporaneous correlation for any pair of 30-minute or 1-day changes in common macroeconomic and financial variables. These high-frequency contemporaneous correlations cannot be that high in practice for two reasons. First, it is well-known that taking first-differences of trending variables reduces their variability and their contemporaneous correlation is smaller than that corresponding to the series in levels. Second, many relationships between macroeconomic and financial variables are in the form of lead-lag which implies that significant portions of dependence between any two variables is not contemporaneous. We verify this empirically by computing the pairwise contemporaneous correlations between all the time series used by Nakamura and Steinsson (2018a) and the main macroeconomic time series available from FRED. Even though these pairwise correlations are typically larger than zero in absolute value, they never exceed 0.5 and they average about 0.2.
@@
To analyze whether the variance of the policy shock is relatively large only for some announcement days and not for others, we compute the sample variance of the policy variable $\mathrm{\widehat{Var}}\left(D_{t}\right)$ over disjoint sub-samples. We consider the full sample from 1/1/2000 to 3/19/2014 and the sample post-2004.242424We consider the sample after 2004 because the data for 2-year forward rates, are available from 2004 onward. The sample sizes are $T=T_{P}+T_{C}=1,236$ $\left(T_{P}=106,\,T_{C}=1130\right)$ and $T=T_{P}+T_{C}=836$ $\left(T_{P}=74,\,T_{C}=762\right)$, respectively. For each sample we construct several sub-samples with different window lengths according to the rule $n_{P}=\left\lfloor T_{P}^{4/5}\right\rfloor$ and $m_{P}=\left\lfloor T_{P}/n_{P}\right\rfloor$ where $n_{P}$ is the number of observations in each window and $m_{P}$ is the number of windows in the treatment sample. [^25] The available data for the control sample is up to and including 2012. Thus, in constructing the sub-samples we consider all announcement and non-announcement days until the end of 2012.262626The choice of the window length $n_{P}=\left\lfloor T_{P}^{4/5}\right\rfloor$ is optimal for nonparametric smoothing under an MSE criterion. The number of windows in the control sample is set equal to that in the treatment sample, i.e., $m_{C}=m_{P}$, and we set $n_{C}=\left\lfloor T_{C}/m_{C}\right\rfloor$ so that each corresponding window in the treatment and control sample brackets the same period.
The evidence in Table 3 shows that there is some time-variation in $\mathrm{\widehat{Var}}\left(D_{t}\right)$ in both the treatment and control sample, though it does not deviate much from the average value computed over the treatment and control sample, respectively. In particular, the ratio of $\mathrm{\widehat{Var}}\left(D_{t}\right)$ in the corresponding treatment and control sub-samples displays some time-variation even though it does not fall substantially below the ratio computed over the full treatment and control sample. For example, for the period 2004-2012 the rule selects two windows. In the first window, the ratio of $\mathrm{\widehat{Var}}\left(D_{t}\right)$ in the treatment and control sample is 26.27. This is only marginally smaller than that corresponding to the sample 2004-2012, 36.72.
@@
Treatment 0.000808 0.001200 Treatment, 1st window 0.000725 0.002011
@@
Ratio 36.72 32.43 Treatment, 2nd window 0.001374 0.001259
@@
As discussed in Section 2.5, an implication of leakage is serial dependence in $D_{t}$ in the days prior to the FOMC announcement. We evaluate this and consider both the 30-minute changes and 1-day changes in the policy variable. For any given announcement day, we estimate the autoregressive coefficients in the regressions,
@@ 4.2 Response of Blue Chip Forecasts on Output to Monetary Policy News
$BCrev_{t}$ $=c+\beta D_{t}+u_{t},$ (4.2)
where $BCrev_{t}$ is the monthly change in Blue Chip survey expectations about real GDP and $D_{t}$ is the policy news shock that occurs in that month. See Bauer and Swanson (2023a) and Nakamura and Steinsson (2018a) for details on how to construct $BCrev_{t}$ and $D_{t}$. It is likely that $D_{t}$ is endogenous for several reasons. As shown by Bauer and Swanson (2023a) $D_{t}$ is correlated with publicly known macroeconomic and financial market data, say $X_{t}$, that predate the FOMC announcement. [^28] The index $t$ of $X_{t}$ should not create confusion. Since these data releases occur before the FOMC announcement, $X_{t}$ collects information that is known before date $t$ but it is still observable at date $t$ or is correlated with variables or news that are realized or occur at time $t$. Since $X_{t}$ is omitted from the regressors, $u_{t}$ and $D_{t}$ are correlated. Nevertheless, the event study regression is valid provided that relative exogeneity holds. The latter requires the policy shock to dominate any other shock in the event window. This means that the variance of $D_{t}$ cannot be an order of magnitude smaller than the variance of $BCrev_{t}.$ $BCrev_{t}$ is constructed as the average of the 1-, 2-, and 3-quarter ahead consensus forecasts. We denote the latter as $BCrev_{t}\mathrm{-1q}$, $BCrev_{t}\mathrm{-2q}$ and $BCrev_{t}\mathrm{-3q}$, respectively.
Table 4 reports the sample variances of $D_{t}$, $BCrev_{t}$, $BCrev_{t}\mathrm{-1q}$, $BCrev_{t}\mathrm{-2q}$ and $BCrev_{t}\mathrm{-3q}$, and of the Treasury yields over the full-sample 1995-2014 as well as over the sub-samples 2000-2014, 2000-2007 and 1995-2000. Strikingly, the variance of $D_{t}$ is much smaller than that of $BCrev_{t}$. For example, in the sample 2000-2014 the variances of $BCrev_{t}$ and $BCrev_{t}\mathrm{-1q}$ are thirteen and thirty four times larger than $\mathrm{\widehat{Var}}(D_{t})$. This is likely due to the fact that the event window for the dependent variable is one month but 30 minutes for the policy variable. Intuitively, while the policy surprise $D_{t}$ is constructed as a 30-minute change, when $Y_{t}$ is the one-month change in the Blue Chip forecasts, it may have too large a variance relative to the policy shock as it aggregates all news and factors that are relevant over the month. [^29] An alternative explanation could be that relative exogeneity holds and $\beta$ is very large in absolute value. We rule out this possibility as $\beta$ would need to be implausibly large in order to generate such a large difference. This implies that relative exogeneity does not provide a good approximation when $Y_{t}$ is a much lower-frequency variable than $D_{t}$, and any correlation between $D_{t}$ and $u_{t}$ is not overwhelmed by the high variance of the policy shock.
@@
This argument should also apply to the event study regression with Treasury yields as dependent variable. However, Bauer and Swanson (2023a) found that in regressions where the dependent variable is the change in an asset price or Treasury yield, controlling for macroeconomic and financial data that predate the FOMC announcement does not change the point estimates and their statistical significance. This difference likely arises because the changes in Treasury yields are constructed using a 30-minute or 1-day window around the announcement so that relative exogeneity is likely to provide a good approximation. As can be seen from Table 4, the changes in Treasury yields based on a 1-day window have an order of magnitude similar to those of the changes in the policy variable. Since relative exogeneity provides a good approximation in this case, controlling for $X_{t}$ does not result in a change in the point estimates even though $D_{t}$ and $X_{t}$ are correlated. The same omitted economic news that predate the announcement do not generate bias when $Y_{t}$ is the change of an asset price or Treasury yield over a similarly-sized narrow window used to construct the policy surprise. The key point here is that the validity of the event study approach does not require the absence of endogeneity. Rather, it requires that the policy shock dominates any other shock that is present in the event window. When this condition holds, identification of causal effects from an event study does not require the inclusion of controls in the event study regression. It is interesting to note that this implies that, when the identification conditions are met, event study regressions are immune to forms of p-hacking that involve searching through different control specifications, a potential strength of the high-frequency event study method.
This discussion suggests that the event study approach is more credible when a narrow window is used for both the dependent and independent variable. Both 30-minute and 1-day windows are good choices as relative exogeneity is more likely to hold. This was also informally discussed on p. 1289 in Nakamura and Steinsson (2018a). If longer windows (e.g., one-month window) are used, as is the case for Blue Chip forecasts, then relative exogeneity is less likely to hold and the researcher should make more effort to appropriately control for omitted variables and reverse causality. For example, orthogonalizing surprises rather than using the original surprises as suggested by Bauer and Swanson (2023a) could provide a solution in these contexts.
@@ 5 Conclusions
We establish nonparametric conditions for identification of casual effects in high-frequency event studies. We show that identification requires a separability condition on the policy shock from the other shocks present in the window, and relative exogeneity which refers to the variance of the policy shock being an order of magnitude larger than that of the other shocks. Under these conditions we establish the causal meaning of the event study estimand, the consistency and asymptotic distribution of the event study estimator and its robustness to nonlinearities. We propose a simple procedure that can be used to assess relative exogeneity as an approximation and we apply it to the study in Nakamura and Steinsson (2018a) on the real effects of the monetary policy.

### v4

@@ 4.1 Response of Interest Rates to Monetary Policy News
For a larger value of $\rho$ it takes larger values of $\delta$ to decrease the MAE and MSE of $\widehat{\beta}_{\mathrm{ES}}$ to those of $\widehat{\beta}_{\mathrm{oracle}}$. For example, for $\rho=0.5$ increasing the variance of the policy variable in the policy sample nine-fold is not sufficient, while increasing it roughly thirty three times is. Similar features are found when the dependent variable is calibrated to the 5-Year real yield.
The bias of $\widehat{\beta}_{\mathrm{oracle}}$ is always smaller than that of $\widehat{\beta}_{\mathrm{ES}}$. This follows because for $\widehat{\beta}_{\mathrm{oracle}}$ positive and negative deviations from the true value $\beta$ cancel each other across simulation replications, while for $\widehat{\beta}_{\mathrm{ES}}$ these deviations tend to be positive since $\rho>0$. On the other hand, these deviations for $\widehat{\beta}_{\mathrm{ES}}$ tend to be smaller in magnitude, making its variance smaller than that of $\widehat{\beta}_{\mathrm{oracle}}$ for large values of $\delta$. Focusing only on bias does not account for the higher precision of $\widehat{\beta}_{\mathrm{ES}}$. A user concerned about both accuracy and precision should rather focus on performance measures such as MSE or MAE, which incorporate both bias and variance.
@@
Values of $\rho=0.50$ and 0.40 represent quite strong empirical contemporaneous correlation for any pair of 30-minute or 1-day changes in common macroeconomic and financial variables. These high-frequency contemporaneous correlations should not be that large in practice for two reasons. First, it is well-known that taking first-differences of trending variables reduces their variability and their contemporaneous correlation is smaller than that corresponding to the series in levels. Second, many relationships between macroeconomic and financial variables are in the form of lead-lag which implies that significant portions of dependence between any two variables is not contemporaneous. We verify this empirically by computing the pairwise contemporaneous correlations between all the time series used by Nakamura and Steinsson (2018a) and the main macroeconomic time series available from FRED. Even though these pairwise correlations are typically larger than zero in absolute value, they never exceed 0.5 and they average about 0.2.
@@
To analyze whether the variance of the policy shock is relatively large only for some announcement days and not for others, we compute the sample variance of the policy variable $\mathrm{\widehat{Var}}\left(D_{t}\right)$ over disjoint sub-samples. We consider the full sample from 1/1/2000 to 3/19/2014 and the sample post-2004.242424We consider the sample after 2004 because the data for 2-year forward rates, are available from 2004 onward. The sample sizes are $T=T_{P}+T_{C}=1,236$ $\left(T_{P}=106,\,T_{C}=1130\right)$ and $T=T_{P}+T_{C}=836$ $\left(T_{P}=74,\,T_{C}=762\right)$, respectively. For each sample we construct several sub-samples with different window lengths according to the rule $n_{P}=\left\lfloor T_{P}^{4/5}\right\rfloor$ and $m_{P}=\left\lfloor T_{P}/n_{P}\right\rfloor$ where $n_{P}$ is the number of observations in each window and $m_{P}$ is the number of windows in the policy sample. [^25] The available data for the control sample is up to and including 2012. Thus, in constructing the sub-samples we consider all announcement and non-announcement days until the end of 2012.262626The choice of the window length $n_{P}=\left\lfloor T_{P}^{4/5}\right\rfloor$ is optimal for nonparametric smoothing under an MSE criterion. The number of windows in the control sample is set equal to that in the policy sample, i.e., $m_{C}=m_{P}$, and we set $n_{C}=\left\lfloor T_{C}/m_{C}\right\rfloor$ so that each corresponding window in the policy and control sample brackets the same period.
The evidence in Table 3 shows that there is some time-variation in $\mathrm{\widehat{Var}}\left(D_{t}\right)$ in both the policy and control sample, though it does not deviate much from the average value computed over the policy and control sample, respectively. In particular, the ratio of $\mathrm{\widehat{Var}}\left(D_{t}\right)$ in the corresponding policy and control sub-samples displays some time-variation even though it does not fall substantially below the ratio computed over the full policy and control sample. For example, for the period 2004-2012 the rule selects two windows. In the first window, the ratio of $\mathrm{\widehat{Var}}\left(D_{t}\right)$ in the policy and control sample is 26.27. This is not much smaller than that corresponding to the sample 2004-2012, 36.72.
@@
Policy 0.000808 0.001200 Policy, 1st window 0.000725 0.002011
@@
Ratio 36.72 32.43 Policy, 2nd window 0.001374 0.001259
@@
As discussed in Section 2.5, an implication of leakage is serial dependence in $D_{t}$ in the days prior to the FOMC announcement. We evaluate this and consider policy variables constructed both as 30-minute and 1-day changes. For any given announcement day, we estimate the autoregressive coefficients in the regressions,
@@ 4.2 Response of Blue Chip Forecasts on Output to Monetary Policy News
$BCrev_{t}$ $=c+\beta D_{t}+\widetilde{u}_{t},$ (4.2)
where $BCrev_{t}$ is the monthly change in Blue Chip survey expectations about real GDP and $D_{t}$ is the policy news shock that occurs in that month. See Bauer and Swanson (2023a) and Nakamura and Steinsson (2018a) for details on how to construct $BCrev_{t}$ and $D_{t}$. It is likely that $D_{t}$ is endogenous for several reasons. As shown by Bauer and Swanson (2023a) $D_{t}$ is correlated with publicly known macroeconomic and financial market data, say $X_{t}$, that predate the FOMC announcement. [^28] The index $t$ of $X_{t}$ should not create confusion. Since these data releases occur before the FOMC announcement, $X_{t}$ collects information that is known before date $t$ but is still observable at date $t$ or is correlated with variables or news that are realized or occur at time $t$. Since $X_{t}$ is omitted from the regressors, $\widetilde{u}_{t}$ and $D_{t}$ are correlated. Nevertheless, the event study regression is valid provided that relative exogeneity holds. The latter requires the policy shock to dominate any other variable in the event window. This means that the variance of $D_{t}$ cannot be an order of magnitude smaller than the variance of $BCrev_{t}.$ $BCrev_{t}$ is constructed as the change in the average of the 1-, 2-, and 3-quarter ahead consensus forecasts. We denote the latter as $BCrev_{t}\mathrm{-1q}$, $BCrev_{t}\mathrm{-2q}$ and $BCrev_{t}\mathrm{-3q}$, respectively.
Table 4 reports the sample variances of $D_{t}$, $BCrev_{t}$, $BCrev_{t}\mathrm{-1q}$, $BCrev_{t}\mathrm{-2q}$ and $BCrev_{t}\mathrm{-3q}$, and of the Treasury yields over the full-sample 1995-2014 as well as over the sub-samples 2000-2014, 2000-2007 and 1995-2000. Strikingly, the variance of $D_{t}$ is much smaller than that of $BCrev_{t}$. For example, in the sample 2000-2014 the variances of $BCrev_{t}$ and $BCrev_{t}\mathrm{-1q}$ are thirteen and thirty four times larger than the variance of $D_{t}$. This is likely due to the fact that the event window for the dependent variable is one month but it is 30 minutes for the policy variable. Intuitively, while the policy surprise $D_{t}$ is constructed as a 30-minute change, when $Y_{t}$ is the one-month change in the Blue Chip forecasts, it may have too large a variance relative to the policy shock as it aggregates all news and factors that are relevant over the month. [^29] An alternative explanation could be that relative exogeneity holds and $\beta$ is very large in absolute value. We rule out this possibility as $\beta$ would need to be implausibly large in order to generate such a large difference. This implies that relative exogeneity does not provide a good approximation when $Y_{t}$ is a much lower-frequency change than $D_{t}$, and any correlation between $D_{t}$ and $\widetilde{u}_{t}$ is not overwhelmed by the high variance of the policy shock.
@@
This argument should also apply to the event study regression with Treasury yields as dependent variable. However, Bauer and Swanson (2023a) found that in regressions where the dependent variable is the change in an asset price or Treasury yield, controlling for macroeconomic and financial data that predate the FOMC announcement does not change the point estimates and their statistical significance. This difference likely arises because the changes in Treasury yields are constructed using a 30-minute or 1-day window around the announcement so that relative exogeneity is likely to provide a good approximation. As can be seen from Table 4, the changes in Treasury yields based on a 1-day window have an order of magnitude similar to those of the policy variable. Since relative exogeneity provides a good approximation in this case, controlling for $X_{t}$ does not result in a change in the point estimates even though $D_{t}$ and $X_{t}$ are correlated. The same omitted economic news that predate the announcement do not generate bias when $Y_{t}$ is the change of an asset price or Treasury yield over a similarly-sized narrow window used to construct the policy surprise. The key point here is that the validity of the event study approach does not require the absence of endogeneity. Rather, it requires that the policy shock dominates any other variable that is present in the event window. When this condition holds, identification of causal effects from an event study does not require the inclusion of controls in the event study regression. It is interesting to note that this implies that, when the identification conditions are met, event study regressions are immune to forms of p-hacking that involve searching through different control specifications, a potential strength of the high-frequency event study method.
This discussion suggests that the event study approach is more credible when a narrow window is used for both the dependent and independent variable. Both 30-minute and 1-day windows are good choices as relative exogeneity is more likely to hold. This was also informally discussed on p. 1289 in Nakamura and Steinsson (2018a). If longer windows (e.g., one-month windows) are used to form the outcome variable, as is the case for Blue Chip forecasts, then relative exogeneity is less likely to hold and the researcher should make more effort to appropriately control for omitted variables and simultaneity. For example, orthogonalizing surprises rather than using the original surprises as suggested by Bauer and Swanson (2023a) could provide a solution in these contexts.
@@ 5 Conclusions
We establish nonparametric conditions for identification of casual effects in high-frequency event studies. We show that identification can be achieved via a separability condition on the policy shock from the other variables present in the window, and relative exogeneity which refers to the variance of the policy shock being an order of magnitude larger than that of the other variables. Under these conditions we establish the causal meaning of the event study estimand, the consistency and asymptotic distribution of the event study estimator and its robustness to nonlinearities. We propose a simple procedure that can be used to assess relative exogeneity as an approximation and apply it to Nakamura and Steinsson’s (2018a) analysis on the real effects of monetary policy.